OpenStudy (anonymous):

How do you find arccot (sqrt(3)/3)?

OpenStudy (anonymous):

\[\cot^{-1} \theta=\frac{\pi}{2}-\tan^{-1} \theta\]

OpenStudy (anonymous):

Would arccot (x) also equal 1/(arctan(x))?

OpenStudy (aravindg):

nope